I need some help, can somebody help me?


(-1/8x^7)(-1/9x)
Thanks

Answer by checkley71(8403)

(-1/8X^7)(-1/9X)=1/72X^8
OR:
(-1/8)X^7*(-1/9)X=X^8/72